OpenStudy (anonymous):

Find the limits, if they exist; i) lim z->0 Izl^2/z ii) lim z->0 (z-re z)/im z

OpenStudy (dape):

First, the definition of a complex limit is: We write \[ \lim_{z->a}f(z)=L, \] If, for every \( \epsilon>0\) there exists a \( \delta > 0\) such that \[ 0<|z-a|<\delta \Rightarrow |f(z)-L|<\epsilon \] Using this, I claim the first limit is 0, since \[ 0<|z-0|<\delta \Rightarrow \left|\frac{|z|^2}{z}-0\right|=|z|<\epsilon \] When \( \epsilon=\delta. \) The second is pretty obvious that it goes to 0. It's easily proved similarly to above.

OpenStudy (dape):

Oops, meant that the second limit goes toward \( i \). Since \[ z=Re(z)+Im(z)i \Rightarrow \frac{z-Re(z)}{Im(z)}=i \] for all z.

OpenStudy (anonymous):

can you also help me out on this question? Let z belongs to C and a,b belong to R be arbitrary. state whether the set is open, closed, or neither i) 1< IzI<2 ii) im z >or=0 and 1<IzI<2

OpenStudy (anonymous):

think i answered this one earlier. maybe someone else asked it

OpenStudy (anonymous):

as for the second one, isn't \[\frac{a+bi-a}{b}=i\]?

OpenStudy (dape):

The first set is open, since all points are interior. The second is neither, since the limit point 1.5 is included, but the limit point 2i is not.

OpenStudy (anonymous):

or because the boundary is included on the real axis but not in the plane

OpenStudy (anonymous):

interior of the plane rather. one thing to be careful of however, is that \(\frac{|z^2|}{z}\neq |z|\) i think

OpenStudy (anonymous):

re(z)=(1/2)(z+z) im(z)=(1/2)(z-z) how does [z-(re(z)]/im(z) =i ?

OpenStudy (anonymous):

\[z=a+bi\] \(re(z)=a, im(z)=b, z-re(z)=a+bi-a=bi\)

OpenStudy (anonymous):

and \(\frac{bi}{b}=i\) i think this is right

OpenStudy (anonymous):

i got it. thanks

OpenStudy (anonymous):

IzI^2/z is not equal to z. because lzl^2=a^2+b^2. what is lzl^2/z =?

OpenStudy (anonymous):

maybe easier if you write \(z=re^{i\theta}\)

OpenStudy (dape):

I never claimed that |z|^2/z = z, I claimed that the absolute value of it was equal: \[ \left| \frac{|z|^2}{z} \right|=\frac{|z|^2}{|z|} = |z| \]

OpenStudy (anonymous):

ah yes, so you did, my fault

OpenStudy (anonymous):

@satellite73 a<arg(z-z state whether the set is open,closed or neither